# UniSite: The First Cross-Structure Dataset and Learning Framework for End-to-End Ligand Binding Site Detection

### v1

@@ 1 Introduction
2) We propose UniSite-1D and UniSite-3D, two end-to-end methods for protein ligand binding site detection. Both models utilize a transformer encoder-decoder architecture, supervised by a set prediction loss with bijective matching. UniSite 1D/3D directly predict $N$ potentially overlapping binding sites without requiring post-processing clustering steps, thus completely resolves Issue 2. The UniSite-1D variant operates exclusively on 1D protein sequence inputs, providing structure-free binding site detection capability. For enhanced performance, the UniSite-3D variant incorporates 3D structural information while maintaining the same end-to-end prediction framework.
3) To overcome the limitations inherent in traditional evaluation methods outlined in Issue 3, we introduce an Average Precision (AP) metric based on Intersection over Union (IoU) for comprehensive binding site assessment. Extensive experiments have demonstrated that the IoU-based AP maintains strong concordance with method rankings under traditional metrics while overcoming their key limitations, providing a more accurate reflection of prediction quality.
@@ 2 UniSite-DS: The First Uniprot-centric Dataset
To construct UniSite-DS, we performed the following search and processing steps: (1) We utilized AHoJ feidakis2022ahoj to systematically search for all protein–ligand interactions in the PDB database berman2003announcing ; (2) To ensure dataset quality, we excluded entries with a resolution greater than 2.5Å or those determined by non-crystallographic methods; (3) Following P2Rank’s filtering criteria p2rank2018 , we removed solvent molecules and ligands composed of fewer than five atoms, resulting in a total of 143,197 protein–ligand interaction entries; (4) For each interaction, binding site residues were identified within a 4.5Å radius of the ligand; (5) We discarded entries with three or fewer binding site residues to eliminate “floating” ligands; (6) Leveraging UniProt’s unique protein sequence identifiers uniprot2019uniprot , we mapped binding site residues from all protein–ligand interactions of each UniProt entry to their corresponding sequences via SIFTS annotations velankar2012sifts , integrating all ligand binding sites across different PDB structures; (7) To eliminate data redundancy among ligand binding sites, we applied Non-Maximum Suppression (NMS) with an Intersection over Minimum (IoM) threshold of 0.7 and an Intersection over Union (IoU) threshold of 0.5, excluding highly overlapping sites. This process resulted in 13,464 distinct UniProt IDs, of which 4,846 contained multiple ligand-binding sites; (8) Based on the criteria from Proteina geffner2025proteina , we set the sequence length threshold to 800; (9) We manually inspected all UniProt IDs with more than ten ligand binding sites, as well as those where a single protein–ligand complex structure contributed three or more binding sites. As a result, we identified 11,510 valid UniProt IDs, including 3,670 with multiple ligand binding sites. The distribution of ligand binding sites is shown in Figure 1.
As the first UniProt-centric dataset, UniSite-DS encompasses 4.81 times more multi-site entries than previous datasets, and covers 2.96 times more UniProt entries than the widely used PDBbind dataset pdbbind2004 , as well as 2.08 times more than sc-PDB scpdb2015 (Figure 1). UniSite-DS eliminates the statistical biases inherent in earlier datasets and significantly expands the available data on multi-site ligand binding sites. Notably, case studies conducted using UniSite-DS (Appendix B) highlighted the limitations of current binding site prediction methods in handling multi-site proteins. This observation motivated us to develop a novel end-to-end method for protein–ligand binding site detection.
@@ 3 The Proposed Methodology
Figure 2: Comparison of detection approaches. (Top) Conventional learning-based binding site detection methods typically employ a discontinuous workflow: first predicting binary masks for residues/atoms, then clustering these masks into distinct binding sites. (Bottom) In contrast, our method directly outputs a set of $N$ potentially overlapping binding sites in a single step.
@@ 3.1 Set prediction loss for binding site detection
UniSite infers a fixed-size set of $N$ predictions $z=\{(p_{i},m_{i})|m_{i}\in\{0,1\}^{L}\}_{i=1}^{N}$ in a single forward pass, where $m_{i}$ represents the predicted binding site, and $p_{i}$ denotes the probability of binding and $\emptyset$ (non-binding) category. Since the ground truth set $|z^{gt}|=N^{gt}$ and the prediction set $|z|=N$ typically have unequal sizes, we assume $N\geq N^{gt}$ and pad the ground truth set with $\emptyset$ (non-binding) tokens. The padded ground truth set is defined as $z^{gt}_{pad}=\{(c_{i}^{gt},m_{i}^{gt})|c_{i}^{gt}\in\{1,\emptyset\},m_{i}^{gt}\in\{0,1\}^{L}\}_{i=1}^{N}$, where $c_{i}^{gt}=1$ indicates a true binding site and $c_{i}^{gt}=\emptyset$ corresponds to the padding.
To train a set prediction model, we require a bijective matching $\sigma$ between the predicted set $z$ and the padded ground truth set $z^{gt}_{pad}$. This bipartite matching is obtained by minimizing the following cost:
[…]
Figure 3: The architecture of UniSite. Our models employ an encoder to extract the residue-level features. Then a decoder module is used to generate embeddings of the $N$ predicted binding sites. Finally, the segmentation module outputs $N$ potentially overlapping binding site predictions. The encoder comprises dual pathways: a sequence encoder and an optional structural encoder, allowing UniSite to operate with either sequence-only input or combined sequence-structure information.
where $\sigma$ is a permutation of $N$ elements and $\mathcal{L}_{\text{match}}$ quantifies the pairwise matching cost between ground truth site $z_{i}^{gt}$ and the prediction with index $\sigma(i)$. Following prior work (detr2020, ; Stewart2016, ), we employ the Hungarian algorithm to compute the optimal matching and use the matching cost defined as:
$\mathcal{L}_{\text{match}}(z_{i}^{gt},z_{\sigma(i)})=-\mathbf{1}_{\{c_{i}^{gt}\neq\emptyset\}}\text{log}p_{\sigma(i)}(c_{i}^{gt})+\mathbf{1}_{\{c_{i}^{gt}\neq\emptyset\}}\mathcal{L}_{\text{mask}}(m_{i}^{gt},m_{\sigma(i)})$
where $\mathcal{L}_{\text{mask}}$ is a binary mask loss, $p_{\sigma(i)}$ and $m_{\sigma(i)}$ denote the predicted probability and the binding site for the $\sigma(i)$-th prediction, respectively. This matching cost considers both the class prediction and the similarity of the predicted and ground truth binding sites. Given the optimal matching $\hat{\sigma}$, we compose a cross-entropy classification loss and a binary mask loss $\mathcal{L}_{\text{mask}}$ for each predicted site to train model parameters:
$\mathcal{L}_{\text{mask\&cls}}(z^{gt},z)=\sum_{i}^{N}{-\log p_{\hat{\sigma}(i)}(c_{i}^{gt})}+\mathbf{1}_{\{c_{i}^{gt}\neq\emptyset\}}\mathcal{L}_{\text{mask}}(m_{i}^{gt},m_{\hat{\sigma}(i)})$
@@ 3.2 UniSite architecture
As illustrated in Figure 3, UniSite comprises three main components: (1) an encoder module that extracts residue-level representations $\mathcal{F}\in\mathbb{R}^{L\times d_{\text{model}}}$ of the protein; (2) a decoder module consisting of multiple Transformer decoder layers to generate embeddings of the $N$ predicted binding sites, and (3) a segmentation module which combines the residue-level representations and the decoder embeddings to produce the final predictions $\{(p_{i},m_{i})|m_{i}\in\{0,1\}^{L}\}_{i=1}^{N}$. This architecture maintains conciseness while demonstrating strong compatibility with existing protein representation methods. In our implementation, we construct two variants: UniSite-1D using only sequence encoding, and UniSite-3D incorporating both sequence and structural encoders.
@@
The amino acid sequence encodes the primary information of a protein and serves as the fundamental input for the UniProt-centric binding site detection. First, an input embedding module generates the initial per-residue features through: (1) learnable embeddings for the 21 amino acid types (20 standard amino acids plus an "unknown" category); (2) sinusoidal positional embedding (attention2017, ) for residue indices; (3) pre-trained ESM-2 (esm2021, ) protein embeddings. These combined features are then processed by a stack of Transformer encoder layers to capture residue-residue interactions and global sequence patterns.
@@
$u_{i}^{(l)}=\text{ReLU}\left(\text{BN}\left(\underset{r\in\mathcal{R}}{\Sigma}W_{r}\underset{j\in\mathcal{N}_{r}(i)}{\Sigma}h_{j}^{(l-1)}\right)\right),\,h_{i}^{(l)}=h_{i}^{(l-1)}+u_{i}^{(l)}$
@@
The decoder comprises multiple Transformer decoder layers (attention2017, ) that simultaneously process $N$ embeddings of dimension $d_{\text{model}}$, $\mathcal{Q}\in\mathbb{R}^{N\times d_{\text{model}}}$, through multi-head self-attention and cross-attention mechanisms. Following established practices in (detr2020, ; maskformer2021, ), these input embeddings are learnable positional embeddings which we refer to as site queries. The attention mechanisms enable the decoder to perform global reasoning over all potential binding sites while incorporating contextual information from the residue-level protein features output by the encoder.
@@
We process the $N$ site queries through a linear classifier followed by the softmax activation to generate class probabilities $\{p_{i}\in\Delta^{2}\}_{i=1}^{N}$. Here, the classifier predicts an additional $\emptyset$ (non-binding) category to indicate when a query does not correspond to any actual binding site. For mask prediction, the site queries $\mathcal{Q}\in\mathbb{R}^{N\times d_{\text{model}}}$ are converted to $N$ mask embeddings $\mathcal{E}_{\text{mask}}\in\mathbb{R}^{N\times d_{\text{model}}}$ by a Multilayer Perceptron (MLP). Finally, the binary mask prediction for each query is computed via dot-production between the $i$-th mask embedding and the residue-level protein features $\mathcal{F}\in\mathbb{R}^{L\times d_{\text{model}}}$, followed by a sigmoid activation:
$m_{i}[j]=\text{sigmoid}\left(\mathcal{E}_{\text{mask}}[i,:]\cdot\mathcal{F}[j,:]^{T}\right)$
@@ 4 Rethinking the Evaluation Metrics for Binding Site Detection
DCC (Distance between the predicted binding site center and the true binding site center) and DCA (Shortest distance between the predicted binding site center and any heavy atom of the ligand) are the two most widely-used metrics for binding site detection. A binding site prediction is considered successful when its DCC/DCA value is below a predetermined threshold. Previous works (p2rank2018, ; deepsite2017, ; deepsurf2021, ; grasp2024, ) quantify prediction performance via the Success Rate, defined as the ratio of successful predictions to the total number of ground truth sites:
@@
However, these metrics suffer from two critical limitations: (Limitation 1) They disregard the prediction scores or ranks, and predictions may be double-counted due to the absence of proper matching criteria (Figure 4 A). (Limitation 2) They only evaluate the center of binding sites and are ligand-dependent (DCC typically considers the ligand center as site center). Since different ligands can bound to one binding site, relying solely on ligand-centered evaluation causes these metrics to completely miss key structural properties such as the shape, size, and residue composition of the binding site. It leads to evaluation failures in certain scenarios (Figure 4 B-D), disregard the crucial information required for downstream tasks (Appendix A).
Previous works have recognized these limitations. To address Limitation 1, a common approach is to calculate the DCC/DCA for either the top-$\bm{n}$ or top-($\bm{n}$+2) predicted binding sites (p2rank2018, ; equipocket2023, ; grasp2024, ), where $\bm{n}$ is the number of ground truth sites. For Limitation 2, DeepSurf (deepsurf2021, ) and Utgés et al. (cmp2024, ) propose computing the IoU between the predicted and ground truth binding site residues. Given two binding sites $m_{A},m_{B}\in\{0,1\}^{L}$, where $L$ is the protein sequence length, the IoU of two sites is defined as:
$\text{IoU}(m_{A},m_{B})=\frac{\text{sum}(m_{A}\&m_{B})}{\text{sum}(m_{A}|m_{B})},\text{where }m_{A},m_{B}\in\{0,1\}^{L}$
@@
To overcome these limitations, we propose to calculate the Average Precision (AP) metric based on the residue-level IoU. We calculate AP as follows: First, we sort all predictions by confidence scores. Then, we match each ground truth site to the predicted site with the highest score and residue-level IoU above a predetermined threshold, enforcing a one-to-one assignment constraint. Finally, we compute AP as the area under the interpolated precision-recall curve following COCO evaluation protocols (coco2014, ), which is widely used in Object Detection. The AP metric offers two significant advantages: (1) the residue-level IoU enables accurate shape and size comparison between binding sites; (2) the one-to-one matching scheme inherently prevents double-counting of predictions.
Figure 4: DCC/DCA failure cases. (A) Repeated counting of the same predicted site since absence of matching. (B) Different ligands bound to the same site lead to deviations in DCC/DCA calculations. (C-D) Failed predictions classified as successful by DCC/DCA but below the IoU threshold.

### v2

@@ 1 Introduction
2) We propose UniSite-1D and UniSite-3D, two end-to-end methods for protein ligand binding site detection. Both models utilize a transformer encoder-decoder architecture, supervised by a set prediction loss with bijective matching. UniSite 1D/3D directly predict NN potentially overlapping binding sites without requiring post-processing clustering steps, thus completely resolves Issue 2. The UniSite-1D variant operates exclusively on 1D protein sequence inputs, providing structure-free binding site detection capability. For enhanced performance, the UniSite-3D variant incorporates 3D structural information while maintaining the same end-to-end prediction framework.
3) To overcome the limitations inherent in traditional evaluation methods outlined in Issue 3, we introduce an Average Precision (AP) metric based on Intersection over Union (IoU) for fair and comprehensive binding site assessment. Extensive experiments have demonstrated that the IoU-based AP maintains strong concordance with method rankings under traditional metrics while overcoming their key limitations, providing a more accurate reflection of prediction quality.
@@ 2 UniSite-DS: The First Uniprot-centric Dataset
To construct UniSite-DS, we performed the following search and processing steps: (1) We utilized AHoJ feidakis2022ahoj to systematically search for all protein–ligand interactions in the PDB database berman2003announcing ; (2) To ensure dataset quality, we excluded entries with a resolution greater than 2.5Å or those determined by non-crystallographic methods; (3) Following P2Rank’s filtering criteria p2rank2018 , we removed solvent molecules and ligands composed of fewer than five atoms, resulting in a total of 143,197 protein–ligand interaction entries; (4) For each interaction, binding site residues were identified within a 4.5Å radius of the ligand; (5) We discarded entries with three or fewer binding site residues to eliminate “floating” ligands; (6) Leveraging UniProt’s unique protein sequence identifiers uniprot2019uniprot , we mapped binding site residues from all protein–ligand interactions of each UniProt entry to their corresponding sequences via SIFTS annotations velankar2012sifts , integrating all ligand binding sites across different PDB structures; (7) To eliminate data redundancy among ligand binding sites, we applied Non-Maximum Suppression (NMS) with an Intersection over Minimum (IoM) threshold of 0.7 and an Intersection over Union (IoU) threshold of 0.5, excluding highly overlapping sites. This process resulted in 13,464 distinct UniProt IDs, of which 4,846 contained multiple ligand-binding sites; (8) Based on the criteria from Proteina geffner2025proteina , we set the sequence length threshold to 800; (9) We manually inspected all UniProt IDs with more than ten ligand binding sites, as well as those where a single protein–ligand complex structure contributed three or more binding sites. As a result, we identified 11,510 valid UniProt IDs, including 3,670 with multiple ligand binding sites. The distribution of ligand binding sites is shown in Figure 1. More details about the UniSite-DS curation workflow and manual inspection process are provided in Appendix B.
As the first UniProt-centric dataset, UniSite-DS encompasses 4.81 times more multi-site entries than previous datasets, and covers 2.96 times more UniProt entries than the widely used PDBbind dataset pdbbind2004 , as well as 2.08 times more than sc-PDB scpdb2015 (Figure 1). UniSite-DS eliminates the statistical biases inherent in earlier datasets and significantly expands the available data on multi-site ligand binding sites. Notably, case studies conducted using UniSite-DS (Appendix E) highlighted the limitations of current binding site prediction methods in handling multi-site proteins. This observation motivated us to develop a novel end-to-end method for protein–ligand binding site detection.
@@ 3 The Proposed Methodology
Figure 2: Comparison of detection approaches. (Top) Conventional learning-based binding site detection methods typically employ a discontinuous workflow: first predicting binary masks for residues/atoms, then clustering these masks into distinct binding sites. (Bottom) In contrast, our method directly outputs a set of NN potentially overlapping binding sites in a single step.
@@ 3.1 Set prediction loss for binding site detection
UniSite infers a fixed-size set of NN predictions $z=\{(p_{i},m_{i})|m_{i}\in\{0,1\}^{L}\}_{i=1}^{N}$ in a single forward pass, where $m_{i}$ represents the predicted binding site, and $p_{i}$ denotes the probability of binding and $\emptyset$ (non-binding) category. Since the ground truth set $|z^{gt}|=N^{gt}$ and the prediction set $|z|=N$ typically have unequal sizes, we assume $N\geq N^{gt}$ and pad the ground truth set with $\emptyset$ (non-binding) tokens. The padded ground truth set is defined as $z^{gt}_{pad}=\{(c_{i}^{gt},m_{i}^{gt})|c_{i}^{gt}\in\{1,\emptyset\},m_{i}^{gt}\in\{0,1\}^{L}\}_{i=1}^{N}$, where $c_{i}^{gt}=1$ indicates a true binding site and $c_{i}^{gt}=\emptyset$ corresponds to the padding.
To train a set prediction model, we require a bijective matching $\sigma$ between the predicted set $z$ and the padded ground truth set $z^{gt}_{pad}$. This matching is obtained by minimizing matching cost $\mathcal{L}_{\text{match}}$:
$\hat{\sigma}=\text{arg min}\,\underset{i}{\overset{N}{\Sigma}}\,\mathcal{L}_{\text{match}}(z_{i}^{gt},z_{\sigma(i)})$ (1)
Figure 3: The architecture of UniSite. Our models employ an encoder to extract the residue-level features. Then a decoder module is used to generate embeddings of the NN predicted binding sites. Finally, the segmentation module outputs NN potentially overlapping binding site predictions. The encoder comprises dual pathways: a sequence encoder and an optional structural encoder, allowing UniSite to operate with either sequence-only input or combined sequence-structure information.
where $\sigma$ is a permutation of NN elements and $\mathcal{L}_{\text{match}}$ quantifies the pairwise matching cost between ground truth site $z_{i}^{gt}$ and the prediction with index $\sigma(i)$. Following prior work (detr2020, ; Stewart2016, ), we employ the Hungarian algorithm to compute the optimal matching and use the matching cost defined as:
$\mathcal{L}_{\text{match}}(z_{i}^{gt},z_{\sigma(i)})=-\mathbf{1}_{\{c_{i}^{gt}\neq\emptyset\}}\text{log}p_{\sigma(i)}(c_{i}^{gt})+\mathbf{1}_{\{c_{i}^{gt}\neq\emptyset\}}\mathcal{L}_{\text{mask}}(m_{i}^{gt},m_{\sigma(i)})$ (2)
where $\mathcal{L}_{\text{mask}}=\lambda_{\text{bce}}\mathcal{L}_{\text{bce}}+\lambda_{\text{dice}}\mathcal{L}_{\text{dice}}$ is a combination of BCE loss and dice loss (diceloss2016, ), $p_{\sigma(i)}$ and $m_{\sigma(i)}$ denote the predicted probability and the binding site for the $\sigma(i)$-th prediction, respectively. This matching cost considers both the class prediction and the similarity of the predicted and ground truth binding sites. Given the optimal matching $\hat{\sigma}$, we compose a cross-entropy classification loss and the binary mask loss $\mathcal{L}_{\text{mask}}$ for each predicted site to train model parameters:
$\mathcal{L}_{\text{mask\&cls}}(z^{gt},z)=\lambda_{\text{cls}}\sum_{i}^{N}{-\log p_{\hat{\sigma}(i)}(c_{i}^{gt})}+\mathbf{1}_{\{c_{i}^{gt}\neq\emptyset\}}\mathcal{L}_{\text{mask}}(m_{i}^{gt},m_{\hat{\sigma}(i)})$ (3)
@@ 3.2 UniSite architecture
As illustrated in Figure 3, UniSite comprises three main components: (1) an encoder module that extracts residue-level representations $\mathcal{F}\in\mathbb{R}^{L\times d_{\text{model}}}$ of the protein; (2) a decoder module consisting of multiple Transformer decoder layers to generate embeddings of the NN predicted binding sites, and (3) a segmentation module which combines the residue-level representations and the decoder embeddings to produce the final predictions $\{(p_{i},m_{i})|m_{i}\in\{0,1\}^{L}\}_{i=1}^{N}$. This architecture maintains conciseness while demonstrating strong compatibility with existing protein representation methods. In our implementation, we construct two variants: UniSite-1D using only sequence encoding, and UniSite-3D incorporating both sequence and structural encoders.
@@
The amino acid sequence encodes the primary information of a protein and serves as the fundamental input for the UniProt-centric binding site detection. First, an input embedding module receives three inputs: (1) learnable embeddings for the 21 amino acid types (20 standard amino acids plus an "unknown" category); (2) sinusoidal positional embedding (attention2017, ) for residue indices; (3) pre-trained ESM-2 (esm2021, ) protein embeddings. These three components are concatenated along the feature dimension, and subsequently processed by a 3-layer multilayer perceptron (MLP) to generate the initial per-residue features. Then a stack of Transformer encoder layers process the combined features to capture residue-residue interactions and global sequence patterns.
@@
$u_{i}^{(l)}=\text{ReLU}\left(\text{BN}\left(\underset{r\in\mathcal{R}}{\Sigma}W_{r}\underset{j\in\mathcal{N}_{r}(i)}{\Sigma}h_{j}^{(l-1)}\right)\right),\,h_{i}^{(l)}=h_{i}^{(l-1)}+u_{i}^{(l)}$ (4)
@@
The decoder comprises multiple Transformer decoder layers (attention2017, ) that simultaneously process NN embeddings of dimension $d_{\text{model}}$, $\mathcal{Q}\in\mathbb{R}^{N\times d_{\text{model}}}$, through multi-head self-attention and cross-attention mechanisms. Following established practices in (detr2020, ; maskformer2021, ), these input embeddings are learnable positional embeddings which we refer to as site queries. The attention mechanisms enable the decoder to perform global reasoning over all potential binding sites while incorporating contextual information from the residue-level protein features output by the encoder.
@@
We process the NN site queries through a linear classifier followed by the softmax activation to generate class probabilities $\{p_{i}=(p_{i}^{site},p_{i}^{\emptyset})|p_{i}^{site},p_{i}^{\emptyset}\in[0,1]\}_{i=1}^{N}$. Here, the classifier predicts an additional $\emptyset$ (non-binding) category to indicate when a query does not correspond to any actual binding site. For mask prediction, the site queries $\mathcal{Q}\in\mathbb{R}^{N\times d_{\text{model}}}$ are converted to NN mask embeddings $\mathcal{E}_{\text{mask}}\in\mathbb{R}^{N\times d_{\text{model}}}$ by a MLP. Finally, the binary mask prediction for each query is computed via dot-production between the $i$-th mask embedding and the residue-level protein features $\mathcal{F}\in\mathbb{R}^{L\times d_{\text{model}}}$, followed by a sigmoid activation:
$m_{i}[j]=\text{sigmoid}\left(\mathcal{E}_{\text{mask}}[i,:]\cdot\mathcal{F}[j,:]^{T}\right)$ (5)
@@ 4 Rethinking the Evaluation Metrics for Binding Site Detection
DCC (Distance between the predicted binding site center and the true binding site center) and DCA (Shortest distance between the predicted binding site center and any heavy atom of the ligand) are the two most widely-used metrics for binding site detection. A binding site prediction is considered successful when its DCC or DCA value is below a predetermined threshold. Previous works (p2rank2018, ; deepsite2017, ; deepsurf2021, ; grasp2024, ) quantify prediction performance via the Success Rate, defined as the ratio of successful predictions to the total number of ground truth sites:
@@
However, these metrics suffer from two critical limitations: (Limitation 1) They disregard the prediction scores or ranks, and predictions may be double-counted due to the absence of proper matching criteria (Figure 4 A and Table S1). (Limitation 2) They only evaluate the center of binding sites and are ligand-dependent (DCC typically considers the ligand center as site center). Since different ligands can bound to one binding site, relying solely on ligand-centered evaluation causes these metrics to completely miss key structural properties such as the shape, size, and residue composition of the binding site. It leads to evaluation failures in certain scenarios (Figure 4 B-D), disregard the crucial information required for downstream tasks (Appendix A).
The quantitative analysis of the DCC and DCA metrics is provided in Appendix G to further substantiate their evaluative flaws. The analysis reveals that approximately 20% of proteins are subject to double-counting during evaluation. Furthermore, the measured mean ground truth DCC (2.15 Å, 92.65% < 4 Å) and DCA (1.57 Å, 98.88% < 4 Å) exhibit a significant deviation from the ideal value of 0, revealing a systematic bias inherent to these metrics. These results directly correspond to the previously discussed limitations and confirm that DCC and DCA metrics significantly distort model performance assessment.
Previous works have recognized these limitations. To address Limitation 1, a common approach is to calculate the DCC or DCA for either the top-$\bm{n}$ or top-($\bm{n}$+2) predicted binding sites (p2rank2018, ; equipocket2023, ; grasp2024, ), where $\bm{n}$ is the number of ground truth sites. For Limitation 2, DeepSurf (deepsurf2021, ) and Utgés et al. (cmp2024, ) propose computing the IoU between the predicted and ground truth binding site residues. Given two binding sites $m_{A},m_{B}\in\{0,1\}^{L}$, where $L$ is the protein sequence length, the IoU of two sites is defined as:
$\text{IoU}(m_{A},m_{B})=\frac{\text{sum}(m_{A}\&m_{B})}{\text{sum}(m_{A}|m_{B})},\text{where }m_{A},m_{B}\in\{0,1\}^{L}$ (7)
@@
To overcome these limitations, we propose to calculate the Average Precision (AP) metric based on the residue-level IoU as a fair metric for method evaluation. We calculate AP as follows: First, we sort all predictions by confidence scores. Then, we match each ground truth site to the predicted site with the highest score and residue-level IoU above a predetermined threshold, enforcing a one-to-one assignment constraint. Finally, we compute AP as the area under the interpolated precision-recall curve following COCO evaluation protocols (coco2014, ), which is widely used in object detection. The pseudo-code for AP calculation is provided in Appendix K. The AP metric offers two significant advantages: (1) the residue-level IoU enables accurate shape and size comparison between binding sites; (2) the one-to-one matching scheme inherently prevents double-counting of predictions.
Figure 4: DCC or DCA failure cases. (A) Repeated counting of the same predicted site since absence of matching. (B) Different ligands bound to the same site lead to deviations in DCC or DCA calculations. (C-D) Failed predictions classified as successful by DCC or DCA but below the IoU threshold.